### v2

# Algebraic Barriers to Halving Algorithmic Information Quantities in Correlated Strings

Abstract
We study the possibility of scaling down algorithmic information quantities in tuples of correlated strings. In particular, we address a question raised by Alexander Shen: whether, for any triple of strings $(a,b,c)$, there exists a string $z$ such that each of the values of conditional Kolmogorov complexity $C(a|z),C(b|z),C(c|z)$ is approximately half of the corresponding unconditional Kolmogorov complexity. We provide a negative answer to this question by constructing a triple $(a,b,c)$ for which no such string $z$ exists. Our construction is based on combinatorial properties of incidences in finite projective planes and relies on recent bounds on point-line incidences over prime fields. As an application, we show that this impossibility implies lower bounds on the communication complexity of secret key agreement protocols in certain settings. These results reveal algebraic obstructions to efficient information exchange and highlight a separation in the information-theoretic behavior of projective planes over fields with and without proper subfields.
1 Introduction
Algorithmic information theory (AIT), introduced and developed in the 1960s by Solomonoff [1, 2, 3], Kolmogorov [4] and Chaitin [5], aims to define the amount of information in a discrete object and to quantify the information shared by several objects. The crucial difference from Shannon’s information theory is that AIT is interested not in an average compression rate (for some distribution of probabilities) but in the optimal compression of some specific individual object. Speaking informally, the amount of information in an individual object (e.g., in a string, in a text, and so on) is defined as the minimal size of a program that produces this specific object. The length of an optimal program producing a string $x$ is called Kolmogorov complexity of $x$ and denoted $C(x)$. Similarly, conditional Kolmogorov complexity of $x$ given $y$, denoted $C(x|y)$, is the length of an optimal program producing a string $x$ given $y$ as an input. A string $x$ is called random or incompressible if $C(x)\approx|x|$.
The value of Kolmogorov complexity depends on the chosen programming language. However, it is known that there exist optimal programming languages that make the complexity function minimal up to bounded additive terms. An extensive introduction to AIT and the theory of Kolmogorov complexity can be found, for example, in the classical paper [6] and in the textbooks [7, 8].
AIT is tightly connected with the classical Shannon’s information theory. The technique of Kolmogorov complexity is used in various problems of theoretical computer science and discrete mathematics. Time bounded Kolmogorov complexity has deep and interesting links with computational complexity and theoretical cryptography, see., e.g., the surveys [9] and [10].
One of the fundamental questions of the AIT is a characterization of the possible values of Kolmogorov complexity of a tuple of strings. For example, for any triple of strings $x,y,z$, we have seven values of Kolmogorov complexity (sometimes called complexity profile of $(x,y,z)$):
$C(x),\ C(y),\ C(z),\ C(x,y),\ C(x,z),\ C(y,z),\ C(x,y,z).$
Which vectors of seven positive numbers can be realized as Kolmogorov complexity of some $x,y,z$? What are the universal constraints connecting different components in these vectors of tuples? There are, for example, classical inequalities
$C(x,y)\leq C(x)+C(y)+O(\log C(x,y))$
(subadditivity, or non-negativity of the mutual information) and
$C(x,y,z)+C(z)\leq C(x,z)+C(y,z)+O(\log C(x,y,z))$
(submodularity, or non-negativity of the conditional mutual information). It is known that for triples of strings there are no substantially different linear inequalities for Kolmogorov complexity: any linear inequality for Kolmogorov complexity of $x,y,z$ (valid up to an additive term $o(C(x,y,z))$) is necessarily a positive linear combination of several instances of subadditivity and submodularity, [11]. However, when four or more strings are involved (so we have $\geq 2^{4}-1=15$ quantities of Kolmogorov complexity), there are some completely different and non-intuitive linear inequality (usually called non Shannon type inequalities) that cannot be represented as linear combinations of subadditivity and submodularity, see e.g. the survey [12]. It is known that exactly the same linear inequalities that are valid for Kolmogorov complexity and for Shannon’s entropy, but the problem of precise characterization of these inequalities for $n\geq 4$ objects remains widely open.
While the questions on linear inequalities for Kolmogorov complexity and for Shannon’s entropy are known to be equivalent, from some other perspectives, the questions on Kolmogorov complexity might look more difficult than similar questions concerning Shannon’s entropy. It is not known, for example, whether the complexity profiles can be scaled with any factor $\lambda>0$ (let us say, up to a logarithmic additive term). More specifically, the following question is open:
Question 1.
Let $\lambda$ be a positive real number. Is it true that for every $k$-tuple of strings $(x_{1},\ldots,x_{k})$ there exists another $k$-tuple $(x_{1}^{\prime},\ldots,x_{k}^{\prime})$ such that
$C(x_{i_{1}}^{\prime},x_{i_{2}}^{\prime},\ldots,x_{i_{s}}^{\prime})=\lambda C(x_{i_{1}},x_{i_{2}},\ldots,x_{i_{s}})+O(\log C(x_{1},\ldots,x_{n}))$
for all tuples of indices $(i_{1},\ldots,i_{s})$, $1\leq i_{1}<i_{2}<\ldots<i_{s}\leq k$
The answer to this question is known to be positive for $k\leq 3$ and any $\lambda$ and for any $k$ and integer $\lambda$. For non-integer factors, e.g., for $\lambda=1/2$, the question is open for all $k\geq 4$, see [13]. Alexander Shen raised another question (see a comment to Question 1 in [13]):
Question 2.
Let $\lambda<1$ be a positive real number. Is it true that for every $k$-tuple of strings $(x_{1},\ldots,x_{k})$ there exists a string $z$ such that
$C(x_{i}\,|\,z)=\lambda C(x_{i})+O(\log C(x_{1},\ldots,x_{n}))$
for $i=1,\ldots,k$
The positive answer to Question 2 would imply the positive answer to Question 1. Besides, Question 2 is interesting in its own right as a special case of the problem of extension of complexity profiles, generalizing the classical notions of common information (by Gács-Körner [14] and Wyner [15]):
Question 3.
(informal) For each given $k$-tuple of strings $(x_{1},\ldots,x_{k})$, what can we say about possible values
$\{C(x_{i_{1}},x_{i_{2}},\ldots,x_{i_{s}},z)\}_{1\leq i_{1}<\ldots<i_{s}\leq k}$
achievable with various strings $z$?
It is known that the answer to Question 2 is positive for $k=1$ and for $k=2$ (see Section 2). The main result of this paper is the negative answer to this question for $k=3$, even for $\lambda=1/2$. In this paper we show there exists a triple of strings $(a,b,c)$ such that there is no $z$ that “halves” the complexities of each of them,
$C(a|z)\approx\frac{1}{2}C(a),\ C(b|z)\approx\frac{1}{2}C(b),\ C(c|z)\approx\frac{1}{2}C(c).$
We provide an example of a triple $(a,b,c)$ such that for every $z$ such that $C(a|z)\approx\frac{1}{2}C(a),\ C(b|z)\approx\frac{1}{2}C(b),$ the value of $C(c|z)$ must be much smaller than $\frac{1}{2}C(c)$. This result is proven in Section 3.
1.1 The main construction
To prove our main result, we propose a fairly explicit construction of a tuple that provides the negative answer to Question 2. This construction is based on incidences in a finite projective plane. We fix a finite field $\mathbb{F}$, take the projective plane over this field, and consider pairs $(x,y)$, where $x$ is a point in this plane and $y$ is a projective line passing through a point. We call such pairs incidences. An incidence in a projective plane is a classical combinatorial object, and their properties were extensively studied in various contexts. Incidences were considered in AIT, see, e.g., [16, 17].
In a projective plane over $\mathbb{F}$ there are $\Theta(|\mathbb{F}|^{2})$ points, $\Theta(|\mathbb{F}|^{2})$ lines, and $\Theta(|\mathbb{F}|^{3})$ incidences. For the vast majority of incidences $(x,y)$ we have
$C(x)\approx 2\log|\mathbb{F}|,\ C(y)\approx 2\log|\mathbb{F}|,C(x,y)\approx 3\log|\mathbb{F}|.$ (1)
The upper bounds are trivial: to specify a point or a line in a projective plane, it is enough to provide two elements of $\mathbb{F}$; to specify together a point and a line incident to this point, it is enough to provide three elements of $\mathbb{F}$. The lower bound follows from a simple counting argument: the number of programs (descriptions) shorter than $k$ is less than $2^{k}$; therefore, for most incidences $(x,y)$ there is no short description, and $C(x,y)\gtrapprox 3\log|\mathbb{F}|$. A similar argument implies $C(x)\gtrapprox 2\log|\mathbb{F}|$ and $C(y)\gtrapprox 2\log|\mathbb{F}|$. We call an incidence $(x,y)$ typical if it satisfies (1).
Thus, for a typical incidence $(x,y)$ the mutual information between $x$ and $y$ is
$I(x:y):=C(x)+C(y)-C(x:y)\approx\log|\mathbb{F}|.$
An. Muchnik observed in [16] that the mutual information of an incidence is hard to “materialize,” i.e., we cannot find a $z$ that “embodies” this amount of information shared by $x$ and $y$. More formally, Muchnik proved that
$\text{there is no }z\text{ such that }C(z|x)\approx 0,\ C(z|y)\approx 0,\ C(z)\approx I(x:y).$ (2)
This insight did not completely resolve the question: the optimal trade-off between $C(z|x)$, $C(z|x)$, $C(z)$ is still not fully understood. Our work follows this direction of research. We prove that for prime fields $\mathbb{F}$, some specific values of $C(z|x)$, $C(z|y)$, and $C(z)$ are forbidden:
[…]
see a more precise statement in Theorem 4 on p. 4. This result contrasts with a much simpler fact proven in [18]:
[…]
(see a more detailed discussion in Section 2).
Our proof of (3) is a fairly direct application of a remarkable result by Sophie Stevens and Frank De Zeeuw, which provides a non-trivial upper bound for the number of incidences between points and lines in a plane over a prime field, [19]. The first theorem of this type was proven by Bourgain, Katz, and Tao, [20]. This result was further improved in [21, 22, 23]. We use the bound from [19], the strongest to date.
Typical incidences in the projective plane over a prime field imply the negative answer to Question 2: if $(x,y)$ is a typical incidence, we let
$a:=x,\ b:=y,\ c:=\langle x,y\rangle$
(where $\langle x,y\rangle$ is the code of the pair $(x,y)$), and prove that for every string $z$ satisfying the conditions $C(a|z)\approx\frac{1}{2}C(a)\text{ that }C(b|z)\approx\frac{1}{2}C(b),$ the value of $C(c|z)$ must be much bigger than $\frac{1}{2}C(c)$, see Theorem 5.
1.2 Application: impossibility results for secret key agreement
The main result (3) can be interpreted as a partial (though very limited in scope) answer to Question 3, as it claims that for some specific pairs $(x,y)$ (typical incidences) there are some limitations for realizable complexity profiles of triples $(x,y,z)$. It is no surprise that this fact can be used to prove certain no-go results in communication complexity, for settings where the participants of the protocol are given such $x$ and $y$ as their inputs. We present an example of such result — a theorem on secret key agreement protocols, as we explain below.
Unconditional secret key agreement is a one of the basic primitive in information-theoretic cryptography, [24]. In the simplest setting, this is a protocol for two parties, Alice and Bob. At the beginning of the communication, Alice and Bob are given some input data, $x$ and $y$ respectively. It is assumed that $x$ and $y$ are strongly correlated, i.e., the mutual information between $x$ and $y$ is non-negligible. Further, Alice and Bob exchange messages over a public channel and obtain (on both sides) some string $w$ that is incompressible (i.e., $C(w)$ is close to its length) and has negligible mutual information with the transcript of the protocol, i.e.,
$C(w\,|\,\text{concatenation of all messages sent by Alice and Bob})\approx|w|.$
Loosely speaking, Alice and Bob transform the mutual information between $x$ and $y$ into a common secret key (which can be later plugged, for example, into a one-time-pad or some other unconditionally secure cryptographic scheme). The secrecy of the key means that an eavesdropper should get (virtually) no information about this key, even having intercepted all communication between Alice and Bob. For a more detailed discussion of the secret key agreement in the framework of AIT we refer the reader to [25, 26].
The challenges in secret key agreement are to (i) maximize the size of the secret key and (ii) to minimize the communication complexity of the protocol (the total length of messages sent to each other by Alice and Bob). It is known that the optimal size of the secret key is equal to the mutual information between $x$ and $y$,
$I(x:y)=C(x)+C(y)-C(x,y).$
(see [25] for the proof in the framework of AIT and [27, 28] for the original result in the classical Shannon’s settings). There exists a communication protocol that allows to produce a secret of optimal size with communication complexity
$\max\{C(x|y),C(y|x)\},$ (5)
see [25], and this communication complexity is tight, at least for some “hard” pairs of inputs $(x,y)$, see [26]. Moreover, some subtler facts are known:
  • the standard protocol achieving (5) (the construction dates back to [27, 28]; see [25] for the AIT version) is highly asymmetric: all messages are sent by only one party (Alice and Bob);
  • for some pairs of inputs $(x,y)$, if we want to agree on a secret key of maximal possible size $I(x:y)$, not only the total communication complexity must be equal to (5), but actually one of the parties (Alice or Bob) must send (5) bits of information, [29];
  • for some pairs of inputs $(x,y)$, the total communication complexity (5) cannot be reduced even if the parties need to agree on a sub-optimal secret key of size $\epsilon n$, see [26].
It remains unknown whether we can always organize a protocol of secret key agreement where the communication complexity (5) is shared evenly by the parties (both Alice and Bob send $\frac{1}{2}C(x|y)$ bits) if they need to agree on a key of sub-optimal size, e.g., $\frac{1}{2}I(x:y)$.
When we claim that communication complexity of a protocol is large in the worst case, i.e., Alice and Bob must send to each other a significant number of bits at least for some pairs of inputs, it is enough to prove this statement of some specific pair of data sets $(x,y)$. Such a proof may become simpler when we use $(x,y)$ with nice combinatorial properties, even though these inputs may look artificial and unusual for practical applications. Such is the case with the mentioned lower bounds for communication complexity proven in [26] and [29]. Both these arguments employ as an instance of a “hard” input $(x,y)$ a typical incidence in a finite projective plane. Thus, t is natural to ask, for these specific $(x,y)$, whether it is possible to agree on a secret key of a sub-optimal key size with a “balanced” communication load (when Alice and Bob send to each other equal number of bits, a half of the total communication complexity for each party). We show, quite surprisingly, that the answer to this question depends on whether the field admits a proper subfield:
Positive result.
If the field $\mathbb{F}_{q}$ contains a subfield of size $\sqrt{q}$, then there exists a balanced communication protocol with communication complexity $\log q$ where
  • Alice sends to Bob $\approx 0.5\log q$ bits,
  • Bob sends to Alice $\approx 0.5\log q$ bits,
and the parties agree on a secret key of length $\approx 0.5\log q$, which is incompressible even conditional on the transcript of the communication between Alice and Bob.
Negative result.
If the field $\mathbb{F}_{q}$ is prime, then in every balanced communication protocol with communication complexity $\log q$ such that
  • Alice sends to Bob $\approx 0.5\log q$ bits,
  • Bob sends to Alice $\approx 0.5\log q$ bits,
the parties cannot agree on a secret key of length $k\gtrapprox\frac{3}{7}\log q$ (the secrecy means that the key must remain incompressible even conditional on the transcript of the communication between Alice and Bob).
For more precise statements see Theorem 7 and Theorem 6 respectively.
1.3 Techniques
A projective plane is a classical geometric object, and combinatorial properties of discrete projective planes have been studied using a wide variety of mathematical techniques. It is no surprise that, in the context of AIT, the information-theoretic properties of incidences in discrete projective planes have been explored through diverse mathematical approaches. In this paper we bring to AIT another (rather recent) mathematical technique that helps distinguish information-theoretic properties of projective planes over prime fields and over fields containing proper subfields.
As we mentioned above, we apply the new approach to the problem of secret key agreement: we consider the setting where Alice and Bob receive as inputs data sets $x$ and $y$ such that $(x,y)$ is a “typical” incidence in a projective plane ($x$ is a line and $y$ is a point incident to this line) over a finite field $\mathbb{F}$ with $n=\lceil\log|\mathbb{F}|\rceil$. We summarize in Table 1 below several technical results concerning this communication problem, and the techniques in the core of these results.
for any protocol of secret key agreement, information-theoretic techniques:
[…]
(not specific for lines and points)
[…]
even for a key of size $\epsilon n$ [26] expander mixing lemma
(applies to all fast mixing graphs)
[…]
the parties agree on a secret key of size $\approx n$, [29] (only assumes that every two points
define a unique line)
for incidences in a plane over a prime field if additive combinatorics, polynomial
[…]
then the size of the secret key $\ll 0.5n$, [this paper] only projective planes over prime fields)
Table 1: Bounds for secret key agreement in the framework of AIT: results and techniques
One of the motivations for writing this paper was to promote the notable results of [20, 21, 22, 23, 19], which presumably can find interesting applications in AIT and communication complexity.
1.4 Organization
The rest of the paper is structured as follows. In Section 2 we briefly discuss (4) (known from [18]). In Section 3 we formally prove our main result (3). In Section 4 we discuss an application of the main result: we show that the performance of the secret key agreement for Alice and Bob given as inputs an incident pair $(x,y)$ (from a projective plane) differs between fields that do and do not contain proper subfields.
1.5 Notation
  • $|\cal S|$ stands for the cardinality of a finite set $\cal S$
  • for a bit string $x$ we denote by $x_{k:m}$ a factor of $x$ that consists of $m-k+1$ bits at the positions between $k$ and $m$ (in particular, $x_{[1:m]}$ is a prefix of $x$ of length $m$);
  • we fix a computable bijection between pairs of binary strings and individual binary strings, and denote by $\langle x,y\rangle$ the binary string assigned to the pair consisting of $x$ and $y$ (we interpret $\langle x,y\rangle$ as a code of the pair $(x,y)$);
  • we denote $\mathbb{FP}$ the projective plane over a finite field $\mathbb{F}$;
  • $G=(R,L;E)$ stands for a bipartite graph where $L\cup R$ (disjoint union) is the set of vertices and $E\subset L\times R$ is the set of edges;
  • $C(x)$ and $C(x|y)$ stand for Kolmogorov complexity of a string $x$ and, respectively, conditional Kolmogorov complexity of $x$ conditional on $y$, see [7, 8]. We use a similar notation for more involved expressions, e.g., $C(x,y)$ stands for $C(\langle x,y\rangle)$ (Kolmogorov complexity of the code representing $x$ and $y$), $C(x,y|v,w)$ stands for $C(\langle x,y\rangle|\langle v,w\rangle)$, i.e., Kolmogorov complexity of the code of the pair consisting of $x$ and $y$ conditional on the code of another pair consisting of $v$ and $w$;
  • we can also talk about Kolmogorov complexity of more complex combinatorial objects (tuples of strings, elements of finite fields, graphs, points and lines in a discrete projective plane, and so on) assuming that each combinatorial object is represented by its code (for some fixed computable encoding rule)
  • $I(x:y):=C(y)+C(y|x)$ and $I(x:y|z):=C(y|z)-C(y|x,z)$ stand for information in $x$ on $y$ and, respectively, information in $x$ on $y$ conditional on $z$
Many natural equalities and inequalities for Kolmogorov complexity are valid only up to a logarithmic additive term, e.g.,
$C(x,y)=C(x)+C(y|x)\pm O(\log n),$
where $n$ is the sum of lengths of $x$ and $y$ (this is the chain rule a.k.a. Kolmogorov–Levin theorem, see [6]). To simplify the notation, we write $A\stackrel{{\scriptstyle\log}}{{\leq}}B$ instead of
$A\leq B+O(\log N),$
where $N$ is the sum of lengths of all strings involved in the expressions $A$ and $B$. Similarly we define $A\stackrel{{\scriptstyle\log}}{{\geq}}B$ (which means $B\stackrel{{\scriptstyle\log}}{{\leq}}A$) and $A\stackrel{{\scriptstyle\log}}{{=}}B$ (which means $A\stackrel{{\scriptstyle\log}}{{\leq}}B$ and $B\stackrel{{\scriptstyle\log}}{{\leq}}A$). For example, the chain rule can be expressed as $C(x,y)\stackrel{{\scriptstyle\log}}{{=}}C(x)+C(y|x);$ the well known fact of symmetry of the mutual information can be expressed as $I(x:y)\stackrel{{\scriptstyle\log}}{{=}}C(x)+C(y)-C(x,y).$
2 Halving complexities of two strings
In this section we discuss the positive answer to Question 2 for $k=1,2$ and $\lambda=1/2$. These results were proven in [18]. Here we recall the main ideas and technical tools behind this argument.
First of all, we observe that Question 2 for $k=1$ and $\lambda=1/2$ is pretty trivial. Given a string $x$ of length $N$, we can try $z=x_{[1:k]}$ for $k=0,\ldots N$. It is clear that for $k=0$ we have $C(x|x_{[1:k]})=C(x)+O(1)$, and for $k=N$ we obtain $C(x|x_{[1:k]})=O(1)$. At the same time, when we add to the condition $z$ one bit, the conditional complexity $C(x|z)$ changes by only $O(1)$. It follows immediately that for some intermediate value of $k$ we obtain $z=x_{[1:k]}$ such that $C(x|z)=\frac{1}{2}C(x)+O(1)$.
This argument employ (in a very naive from) the same intuition as the intermediate value theorem for continuous functions. The case $k=2$ is more involved, but it also can be proven with “topological” considerations.
Theorem 1.
For all strings $a,b$ of complexity at most $n$ there exists a string $z$ such that
$\left|C(a|z)-\frac{1}{2}C(a)\right|=O(\log n)\text{ and }\left|C(b|z)-\frac{1}{2}C(b)\right|=O(\log n).$
In fact, [18] proved a tighter and more general statement:
Theorem 2.
[18, Theorem 4] For some constant $c$ the following statement holds: for every two strings $a,b$ of complexity at most $n$ and for every integers $\alpha,\beta$ such that
  • $\alpha\leq C(a)-c\log n$,
  • $\beta\leq C(b)-c\log n$,
  • $-C(a|b)+c\log n\leq\beta-\alpha\leq C(b|a)-c\log n$,
there exists a string y such that $|C(a|z)-\alpha|\leq c\text{ and }|C(b|z)-\beta|\leq c.$
With $\alpha=\frac{1}{2}C(a)$ and $\beta=\frac{1}{2}C(b)$, this theorem implies the following corollary, which is (for non-degenerate parameters) a stronger version of Theorem 1:
Corollary 1.
For some constant $c$ the following statement holds: for every two strings $a,b$ such that
$C(a|b)\gg\log n\text{ and }C(b|a)\gg\log n$
there exists a string $z$ such that
$\left|C(a|z)-\frac{1}{2}C(a)\right|\leq c\text{ and }\left|C(b|z)-\frac{1}{2}C(b)\right|\leq c.$
The proof of Theorems 2 proposed by A. Shen involves topological ideas. The argument in a nutshell: we build $z$ by combining two parts, a piece extracted from $a$ and a piece extracted from $b$; the only challenge is to choose the sizes of these two parts in a suitable way. It turns out that suitable sizes of these pieces can be chosen using the topological statement known as the drum theorem, [30], which is equivalent to the fact that a circle is not a retract of a closed disk, see Appendix for more details.
3 Typical incidences in a projective plane
[…]
3.1 Typical pairs: interface between the combinatorial language and AIT
In this section we introduce the framework that helps to translated information-theoretic questions in the combinatorial language.
Definition 1.
Let $G=(L,R;E)$ with $E\subset L\times R$ be a simple non-directed bipartite graph. This graph is bi-regular if all vertices in $L$ have the same degree (the same number of neighbors in $L$) and all vertices in $R$ have the same degree (the same number of neighbors in $L$).
To specify the quantitative characteristics of $G$ we will use a triple of parameters $(\alpha,\beta,\gamma)$ such that
$|L|=2^{\alpha+\gamma},\ |R|=2^{\beta+\gamma},\ |E|=2^{\alpha+\beta+\gamma}.$
If $G$ is bi-regular, then the degrees of vertices in $L$ are equal to $2^{\beta}$ and the degrees of vertices in $R$ are equal to $2^{\alpha}$.
Proposition 1.
Let $G=(L,R;E)$ be a bi-regular with parameters $(\alpha,\beta,\gamma)$, as defined above. If the graph is given explicitly (the complete list of vertices and edges of the graph can be found algorithmically given the parameters $n$ and $d$) the vast majority (let us say, for $99\%$) of pairs $(x,y)\in E$ we have
$\left\{\begin{array}[]{l}C(x)\stackrel{{\scriptstyle\log}}{{=}}\alpha+\gamma,\\ C(y)\stackrel{{\scriptstyle\log}}{{=}}\beta+\gamma,\\ C(x,y)\stackrel{{\scriptstyle\log}}{{=}}\alpha+\beta+\gamma.\end{array}\right.$ (6)
Proof.
This proposition follows from a standard counting, see e.g. [8]. ∎
Definition 2.
For a graph $G=(L,R;E)$ with parameters $(\alpha,\beta,\gamma)$ we say that an edge $(u,v)\in E$ is typical if it satisfies (6).
Proposition 2.
Let $G=(L,R;E)$ be an explicitly given bi-regular bipartite graph with parameters $(\alpha,\beta,\gamma)$. Let $(x,y)\in E$ be a typical edge in this graph. Let be a string $z$ such that
$C(x|y,z)=\alpha,\ C(y|x,z)=\beta,\ I(x:y|z)=\gamma.$
Then there exist an induced subgraph $H=(L^{\prime},R^{\prime};E^{\prime})$ in $G$,
$E^{\prime}\subset L,\ R^{\prime}\subset R,\ E^{\prime}=(L^{\prime}\times R^{\prime})\cap E$
such that
$|L^{\prime}|=2^{\alpha+\gamma+O(\log n)},\ |R^{\prime}|=2^{\beta+\gamma+O(\log n)},\ |E^{\prime}|\geq 2^{\alpha+\beta+\gamma-O(\log n)}.$
Sketch of the proof.
We let
$L^{\prime}=\{x^{\prime}\in L\ :\ C(x^{\prime}|z)\leq C(x|z)\},\ R^{\prime}=\{y^{\prime}\in R\ :\ C(y^{\prime}|z)\leq C(y|z)\}.$
Observe that $x\in L^{\prime}$ and $y\in R^{\prime}$.
Lemma 1.
$|L^{\prime}|=2^{\alpha+\gamma\pm O(\log n)},\ |R^{\prime}|=2^{\beta+\gamma\pm O(\log n)}.$
Proof of lemma.
This lemma is a pretty standard translation between the combinatorial and the information-theoretic languages, see, e.g., [29, lemma 1 and lemma 2] for the upper and lower bound respectively. We sketch the proof here.
Upper bound: For every $x^{\prime}\in L^{\prime}$ there exists at least one program of length $\leq C(x|z)$ that takes $z$ as input and produces $x^{\prime}$. The total number of such programs (binary strings of length $\leq C(x|z)$) is less than $2^{C(x|z)+1}$. It follows that
$|L^{\prime}|\leq 2^{C(x|z)+1}=2^{\alpha+\gamma+O(\log n)}.$
The proof of the upper bound for $R^{\prime}$ is similar.
Lower bound: Given $z$ and the binary expansion of the number $C(x|z)$, we can run in parallel all programs of length $\leq C(x|z)$ feeding them string $z$ as input. Over time, some of these programs will print their results and stop. Thus, we will reveal one by one the elements $x^{\prime}\in L^{\prime}$. The original string $x$ will appear in this enumeration sooner or later. Therefore, to specify $x$ given $z$, we need (i) the information required to organize this enumeration (basically, we need to know the binary expansion of $C(x|z)$, which requires only $O(\log n)$ bits), and (ii) the ordinal number of $x$ in this enumeration. The latter requires no more than $\log|L^{\prime}|$ bits. Hence,
$C(x|z)\leq\log|L^{\prime}|+O(\log n).$
This imply the desired lower bound $|L^{\prime}|\geq 2^{C(x|z)-O(\log n)}=2^{\alpha+\gamma-O(\log n)}$. The proof of the lower bound for $R^{\prime}$ is is absolutely analogous. ∎
It remains to prove a bound on the cardinality of $E^{\prime}$. Given a string $z$, we can run in parallel all programs of length $\alpha+\gamma$ and $\beta+\gamma$ on input $z$ and enumerate the results that they produce. These result will provide us with the lists of elements $L^{\prime}$ and $R^{\prime}$ revealing step by step. Accordingly, we can enumerate edges of $E^{\prime}$. Every pair $(x^{\prime},y^{\prime})\in E^{\prime}$ can be specified by (i) the binary expansion of the numbers $\alpha,\beta,\gamma$ and (ii) by the ordinal number of $(x^{\prime},y^{\prime})$ in the enumeration of $E^{\prime}$. This argument applies in particular to the pair $(x,y)$, which certainly belongs to $E^{\prime}$. Therefore,
$C(x,y|z)\leq\log|E^{\prime}|+O(\log n).$
Reading this inequality from the right to the left, we obtain
$|E^{\prime}|\geq 2^{C(x,y|z)-O(\log n)}=2^{\alpha+\beta+\gamma-O(\log n)},$
and we are done. ∎
3.2 Typical incidences in a projective plane
Now we instantiate the framework discussed above and discuss the central construction of this paper — typical incidences in a finite projective plane.
Example 1.
Let $\mathbb{F}$ be a finite field and $\mathbb{FP}$ be the projective plane over this field. Let $L$ be the set of points and $R$ be the set of line in this plane. A pair $(x,y)\in L\times R$ is connected by an edge iff the chosen point $x$ lies in in the chosen line $y$. Hereafter we denote this graph by $G^{\mathrm{PL}}_{\mathbb{F}}$.
[…]
Theorem 3 (see [17]).
[…]
$C(x)\stackrel{{\scriptstyle\log}}{{=}}2n,\ C(y)\stackrel{{\scriptstyle\log}}{{=}}2n,\ C(x,y)\stackrel{{\scriptstyle\log}}{{=}}3n,$
and there exists a $z$ such that
$C(x|z)\stackrel{{\scriptstyle\log}}{{=}}n,\ C(y|z)\stackrel{{\scriptstyle\log}}{{=}}n,\ C(x,y|z)\stackrel{{\scriptstyle\log}}{{=}}1.5n$
or, equivalently
$C(x|y,z)\stackrel{{\scriptstyle\log}}{{=}}0.5n,\ C(y|x,z)\stackrel{{\scriptstyle\log}}{{=}}0.5n,\ I(x:y|z)\stackrel{{\scriptstyle\log}}{{=}}0.5n,$
as shown in the diagram in Fig. 1.
$n/2$$n/2$$n/2$$n/2$$n/2$$n/2$$x$$y$$z$
Figure 1: Complexity profile for $(x,y,z)$ from Theorem 3.
Sketch of the proof.
The first claim of the theorem (the values of unconditional Kolmogorov complexity) follows from Proposition 1 and from typicality of $(x,y)$. The second claim, concerning the values of conditional Kolmogorov complexity, is more subtle and requires a construction. The statement statement boils down to the following combinatorial claim: the graph $G^{\mathrm{PL}}_{\mathbb{F}}$ can be covered by a relatively small family [^1] More technically, we need $2^{1.5n+O(\log n)}$ such subgraphs, and every edge of $G^{\mathrm{PL}}_{\mathbb{F}}$ is covered by at most $\mathrm{poly}(n)$ subgraphs $H_{i}$. of induced subgraphs $H_{i}=(L_{i},R_{i};E_{i})$, where
$|L_{i}|=|R_{i}|=2^{n},\ |E_{i}|=2^{1.5n}.$
This combinatorial claim, in turn, follows from the facts that $G^{\mathrm{PL}}_{\mathbb{F}}$ is edge-transitive and $\mathbb{F}$ contains a subfield of size $\sqrt{|\mathbb{F}|}$, see [17] for details. ∎
Theorem 3 contrasts with Theorem 4.
Theorem 4.
[…]
$C(x)\stackrel{{\scriptstyle\log}}{{=}}2n,\ C(y)\stackrel{{\scriptstyle\log}}{{=}}2n,\ C(x,y)\stackrel{{\scriptstyle\log}}{{=}}3n,$
and for every $z$ such that
$C(x|y,z)\stackrel{{\scriptstyle\log}}{{=}}(0.5-\epsilon)n,\ C(y|x,z)\stackrel{{\scriptstyle\log}}{{=}}(0.5-\epsilon)n$ (7)
we have $I(x:y|z)\stackrel{{\scriptstyle\log}}{{\geq}}\frac{4}{7}n-O(\epsilon n)\gg 0.5n$.
Proof.
Again, the first claim of the theorem (the values of unconditional Kolmogorov complexity) follows from Proposition 1 and from typicality of $(x,y)$. We proceed with the second claim and assume that $z$ is a string satisfying (7). We denote $\ell=I(x:y|z)$. From Proposition 2 it follows that in $G^{\mathrm{PL}}_{\mathbb{F}}$ there is a subgraph $G^{\prime}=(L^{\prime},R^{\prime},E^{\prime})$ such that
$\begin{array}[]{rcl}|L^{\prime}|&=&2^{C(x|z)+O(\log n)}=2^{(0.5-\epsilon)n+\ell+O(\log n)}\\ |R^{\prime}|&=&2^{C(y|z)+O(\log n)}=2^{(0.5n-\epsilon)+\ell+O(\log n)}\end{array}$ (8)
and
$|E^{\prime}|\geq 2^{C(x,y|z)-O(\log n)}=2^{n-O(\epsilon n)+\ell-O(\log n)}.$ (9)
On the other hand, it was shown in [19] that for every subgraph $G^{\prime}$ in $G^{\mathrm{PL}}_{\mathbb{F}}$ for a prime $\mathbb{F}$ satisfying the constraints
$|L^{\prime}|^{7/8}<|R^{\prime}|<|L^{\prime}|^{8/7}\text{ and }\max\{|L^{\prime}|,|R^{\prime}|\}\leq|\mathbb{F}|^{8/7}$
we have
$|E^{\prime}|\leq(|L^{\prime}|\cdot|R^{\prime}|)^{11/15}.$
We plug in this inequality (8) and (9) and obtain
$n+\ell\leq\frac{22}{15}(0.5n+\ell)+O(\epsilon n)+O(\log n),$
which implies $\ell\geq 4n/7-O(\epsilon n)-O(\log n)$, and we are done. ∎
Theorem 5.
For every $n$ there exists a triple of strings $(a,b,c)$ such that there is no $z$ such that
$\begin{array}[]{rcl}C(a|z)&\stackrel{{\scriptstyle\log}}{{=}}&\frac{1}{2}C(a),\\ C(b|z)&\stackrel{{\scriptstyle\log}}{{=}}&\frac{1}{2}C(b),\\ C(c|z)&\stackrel{{\scriptstyle\log}}{{=}}&\frac{1}{2}C(c).\end{array}$
Proof.
We fix an integer $n$ and the minimal prime number $p$ such that $2^{n}<p<2^{n+1}$, and take a pair $(x,y)$ from Theorem 4 (i.e., a typical edge in $G^{\mathrm{PL}}_{\mathbb{F}_{p}}$). We let
$a:=x,\ b:=y,\ c:=\langle x,y\rangle,$
i.e., $c$ is the code of the pairs $(x,y)$. Then we apply Theorem 4, and we are done. ∎
4 Secret key agreement
In this section we study communication complexity of the protocol of unconditional (information-theoretic) secret key agreement. Let is recall the settings of the unconditional secret key agreement. We deal with two parties, Alice and Bob. Alice and Bob receive input data, $x$ and $y$ respectively. It is assumed that the mutual information between $x$ and $y$ is non-negligible, and its value is known to Alice and Bob, as well as to the adversary. Further, Alice and Bob exchange messages over a public channel and obtain (on both sides) some string $w$ that must be incompressible (i.e., $C(w)$ is close to its length) and must have negligible mutual information with the transcript of the protocol, i.e.,
$C(w\,|\,\text{concatenation of all messages sent by Alice and Bob})\approx|w|.$
Thus, Alice and Bob use the mutual information between $x$ and $y$ to produce a common secret key $w$ using a communication via a non-protected channel. The protocol succeed if Alice and Bob obtain one and the same $w$, and an eavesdropper gets only negligible information about this key, even having intercepted all messages sent to each other by Alice and Bob. In this paper we assume that the communication protocols are deterministic. All arguments easily extends to randomized communication protocols with a public [^2] The case of private sources of randomness is a more complex setting. We leave the consideration of this type of protocols for further research. source of randomness (accessible to Alice, Bob, and Charlie). A more detailed discussion of the settings of secret key agreement problem in the framework of AIT can be found in [25, 26].
The optimal size of the secret key is known to be equal to the mutual information between $x$ and $y$, and communication complexity of the protocol is at most (5), see [25] (in what follows we discuss pairs $(x,y)$ with a symmetric complexity profile where $C(x|y)=C(y|x)$).
4.1 Specific input data: secret key agreement with a typical incidence from a finite plane
Let us focus on the case where the inputs $(x,y)$ represent a pair of typical incidences in a projective plane over a finite field $\mathbb{F}$ (we denote $n:=\lceil\log|\mathbb{F}|\rceil$). In this case the upper bound (5) (which rewrites in this case to to $n$) is tight, the communication complexity cannot be made better than $n-O(\log n)$, [26]. Moreover,
  • (i)
    for every communication protocol, for its transcript $t$ we have
    $I(t:x|y)+I(y:t|x)\stackrel{{\scriptstyle\log}}{{\geq}}n,$ (10)
    see [26] (it is easy to verify that $I(t;x|y)+I(y:t|x)$ is a lower bound for the communication complexity);
  • (ii)
    this bound remains valid even if the parties agree on a sub-optimal secret key of size $\epsilon n$,[26];
  • (iii)
    if Alice and Bob agree on a secret key $w$ of maximal possible size $I(x:y)=n$, then not only the total communication complexity must be equal to $n$ but actually one of the parties (Alice or Bob) must send (5) bits of information, [29].
We summarize:
  • even for a suboptimal key size communication complexity of the protocol $\stackrel{{\scriptstyle\log}}{{\geq}}n$;
  • for an optimal key size the communication is very asymmetric—all $n$ bits are sent by one of the participants.
There remained a question: Does there exist a protocol with a symmetric communication load (both Alice and Bob send $\approx n/2$ bits) with a suboptimal key size? In what follows we show that the answer to this question depends on whether the underlying field contains a proper subfield.
4.2 Prime field: a negative result
Theorem 6.
Let $q$ be a prime number and $\mathbb{F}_{q}$ be the field with $q$ elements. Let $\mathbb{FP}$ be the projective plane over $\mathbb{F}_{q}$ , and $(x,y)$ be a typical incidence in this plane ($x$ is a line in this projective plane, $y$ is a point in this line, and $C(x,y)\stackrel{{\scriptstyle\log}}{{=}}3\log q$). We consider communication protocols where Alice is given as her input $x$ and Bob is given as his input $y$. Assume that such a protocol if communicationally balanced protocol where
  • Alice sends a message $m_{A}$ of length $(\frac{1}{2}+\epsilon)n$ bits to Bob,
  • Bob sends a message $m_{B}$ of length $(\frac{1}{2}+\epsilon)n$ bits to Alice,
  • then Alice and Bob agree on a secret key $w$ of length $k$, satisfying
    $C(w|\langle m_{A},m_{B}\rangle)\stackrel{{\scriptstyle\log}}{{\geq}}k,$
    where $n=\lceil\log q\rceil$, i.e., the protocol reveals virtually no information about the secret to the eavesdropper.
Then $k\stackrel{{\scriptstyle\log}}{{\leq}}3n/7+O(\epsilon n).$ Thus, for small enough $\epsilon$, the size of the secret key $k\ll n/2$.
$(0.5-\epsilon)n$$(0.5-\epsilon)n$$k$$n-k$$(0.5+\epsilon)n$$(0.5+\epsilon)n$$x$$y$$z$
Figure 2: Complexity profile for $(x,y,z)$ from Theorem 6.
Proof.
Let Alice and Bob agree on a secret key $w$ in protocol with transcript $t$. The fact that both Alice and Bob compute $w$ at the end of the protocol means that $C(w|t,x)$ and $C(y|t,y)$ are negligibly small. Security of the key means that $I(w:t)$ is negligible, i.e., the transcript divulges virtually no information about the key. Having these observations, we define $z=\langle t,w\rangle$. We have
$C(z|x,y)=O(\log n)$
(given both $x$ and $y$, we can simulate the protocol and compute the transcript and the key). It follows that $C(t|x)\stackrel{{\scriptstyle\log}}{{=}}I(t:y|x)$ and $C(t|y)\stackrel{{\scriptstyle\log}}{{=}}I(t:x|y)$.
We will assume without loss of generality that
$\begin{array}[]{rcl}C(t|x)&\stackrel{{\scriptstyle\log}}{{=}}&(0.5+\epsilon)n,\\ C(t|y)&\stackrel{{\scriptstyle\log}}{{=}}&(0.5+\epsilon)n,\\ C(t)&\stackrel{{\scriptstyle\log}}{{=}}&(1+2\epsilon)n.\end{array}$
This implies
$C(x|t,y)\stackrel{{\scriptstyle\log}}{{=}}(0.5-\epsilon)n\text{ and }C(y|t,x)\stackrel{{\scriptstyle\log}}{{=}}(0.5+\epsilon)n.$
Finally, Kolmogorov complexity of $z=\langle t,w\rangle$ is equal to $C(t)+C(w)$ (since the mutual information between $w$ and $t$ is negligible); however, conditional on $x$ and conditional on $y$, Kolmogorov complexities of $z$ and $t$ are essentially the same (given the transcript $t$ and the input of one of the parties, we can obtain the secret key $w$ for free). It follows that
$I(x:y|z)\stackrel{{\scriptstyle\log}}{{=}}I(x:y|t)-C(w)\stackrel{{\scriptstyle\log}}{{\leq}}n-k,$
see Fig. 2. Now we apply Theorem 4 and obtain
$n-k\stackrel{{\scriptstyle\log}}{{\geq}}4n/7-O(\epsilon n).$
4.3 Field with a large subfield: a positive result
Theorem 7.
Let $\mathbb{F}_{q}$ be a field with $q$ elements, and $q=p^{2}$ for some integer $p$ (e.g., $p$ is prime and $q$ is a square of this prime number, or $p=2^{k}$ and $q=2^{2k}$).
Let $\mathbb{FP}$ be the projective plane over $\mathbb{F}_{q}$ , and $(x,y)$ be a typical incidence in this plane ($x$ is a line in this projective plane, $y$ is a point in this line, and $C(x,y)=3\log q\pm O(\log n)$). We consider communication protocols where Alice is given as her input $x$ and Bob is given as his input $y$. We claim that there exists a communicationally balanced protocol where
  • Alice sends a message $m_{A}$ of length $n/2$ bits to Bob,
  • Bob sends a message $m_{B}$ of length $n/2$ bits to Alice,
  • Then Alice and Bob compute a secret key $w$ of length $n/2$ such that
    $C(w|\langle m_{A},m_{B}\rangle)\stackrel{{\scriptstyle\log}}{{\geq}}n/2,$
    where $n=\lceil\log q\rceil$, i.e., the protocol reveals virtually no information about the secret to the eavesdropper.
Proof.
A point $x$ and a line $y$ in the projective plane $\mathbb{FP}$ can be specified by their projective coordinates $(x_{0}:x_{1}:x_{2})$ and $(y_{1}:y_{2}:y_{3})$ respectively. Without loss of generality, we will set $x_{0}\not=0$ and $y_{2}\not=0$ and denote
$x_{1}^{\prime}:=x_{1}/x_{0},\ x_{2}^{\prime}:=-x_{2}/x_{0}\text{ and }y_{0}^{\prime}:=y_{0}/y_{2},\ y_{1}^{\prime}:=y_{1}/y_{2}.$
The incidence of $x$ and $y$ means that $x_{0}y_{0}+x_{1}y_{1}+x_{2}y_{2}=0,$ or equivalently
$y_{0}^{\prime}+x_{1}^{\prime}y_{1}^{\prime}-x_{2}^{\prime}=0.$ (11)
Since $q=p^{2}$, the field $\mathbb{F}_{q}$ contains a subfield $\mathbb{G}$ of size $p$, and there exists an element $\xi\in\mathbb{F}_{q}$ such that every element $\alpha\in\mathbb{F}_{q}$ can be represented as
$\alpha=a_{0}+a_{1}\cdot\xi$
for some $a_{0},a_{1}\in\mathbb{G}$. So we may represent $x_{i}^{\prime}$ and $y_{i}^{\prime}$ as follows:
$x^{\prime}_{1}=f+r\xi,\ y^{\prime}_{1}=h+s\xi,\ y^{\prime}_{0}=g+t\xi$
for some $f,g,h,r,s,t\in\mathbb{G}$. In this notation, (11) rewrites to
$(g+t\xi)+(f+r\xi)(h+s\xi)=x_{2}^{\prime}.$ (12)
It follows that
$x_{2}^{\prime}=g+fh+(t+fs+hr)\xi+rs\xi^{2}$
(The value $\xi^{2}$ can be represented as $u+v\xi$ for some $u,v\in\mathbb{G}$, but we do not need to specify these parameters.) Let us recall that Alice knows all parameters derived from $x$ (including $f,r,x_{2}^{\prime}$), and Bob knows all parameters derived from $y$ (including $g,h,s,t$).
Communication protocol.
Round 1: Bob sends to Alice the value $m_{1}:=s$ (this message consists of $\log|\mathbb{G}|=n/2$ bits of information)
Round 2: Alice computes $m_{2}=g+fh$ and sends it to Bob (this message also consists of $\log|\mathbb{G}|=n/2$ bits of information)
Post-processing: Both participants compute the value $f$ and take it as the final result (the secret key, which also consists of $\log|\mathbb{G}|=n/2$ bits of information).
Claim 1: Alice has enough information to compute $m_{2}$.
Proof of claim.
Initially, Alice is given the values of $x_{1}^{\prime}=f+r\xi$ and $x_{2}^{\prime}=u^{\prime}+v^{\prime}\xi$, where $f,r,u^{\prime},v^{\prime}$ are elements of $\mathbb{G}$. When she receives from Bob $s$, she gets all information to compute $rs\xi^{2}=u^{\prime\prime}+v^{\prime\prime}\xi$ (for some $u^{\prime\prime},v^{\prime\prime}\in\mathbb{G}$). From (12) it follows that $g+fh=u^{\prime}-u^{\prime\prime}.$
Alice is given the secret key $f$ as a part of her input. Bob, however, needs to do some computation to get this value.
Claim 2: Bob has enough information to compute the final result $f$.
Proof of claim.
Initially, Bob is given the values $h,s,g,t\in\mathbb{G}$. Bob receives from Alice the value $g+fh,$ which is another element of the field $\mathbb{G}$. This allows him to compute $f$ as $((g+fh)-g)\cdot h^{-1}.$
It remains to show that we reveal no information to the eavesdropper. The adversary can intercept the messages $m_{1}=s$ and $m_{2}=g+fh$. We need to show that these messages give no information about the produced secret key:
Claim 3: $I(f:\langle m_{1},m_{2}\rangle)=O(\log n)$.
Proof of claim.
To specify the incidence $(x,y)$, it is enough to provide the values $f,g,h,r,s,t\in\mathbb{G}$. Thus, we have
$\begin{array}[]{rcl}C(x,y)&\stackrel{{\scriptstyle\log}}{{=}}&C(f,g,h,s,r,t)\\ &\stackrel{{\scriptstyle\log}}{{\leq}}&C(m_{1})+C(m_{2})+C(f,g,h,s,r,t|m_{1},m_{2})\\ &\stackrel{{\scriptstyle\log}}{{\leq}}&C(m_{1})+C(m_{2})+C(s|m_{1},m_{2})+C(f|m_{1},m_{2})+C(h)\\ &&{}+C(g|m_{1},m_{2},f,h)+C(r)+C(t)\\ &\stackrel{{\scriptstyle\log}}{{\leq}}&C(m_{1})+C(m_{2})+C(h)+C(f|m_{1},m_{2})+C(r)+C(t)\\ &\stackrel{{\scriptstyle\log}}{{\leq}}&5\log|\mathbb{G}|+C(f|m_{1},m_{2})\\ &\stackrel{{\scriptstyle\log}}{{=}}&5n/2+C(f|m_{1},m_{2})\end{array}$
(in this calculation, $C(s|m_{1},m_{2})$ vanishes since $m_{1}=s$, and $C(g|m_{1},m_{2},f,h)$ vanishes since we can compute $g$ given $f,h$ and the value of $g+fh$).
Since the incidence $(x,y)$ is typical , i.e., $C(x,y)\stackrel{{\scriptstyle\log}}{{=}}3n$, we obtain $C(f|m_{1},m_{2})\stackrel{{\scriptstyle\log}}{{\geq}}n/2$. Thus, $C(f|m_{1},m_{2})\stackrel{{\scriptstyle\log}}{{=}}C(f)$, and the claim is proven. ∎
5 Conclusion
We have shown that the Kolmogorov complexities of a triple of correlated strings cannot, in general, be reduced by a constant factor through conditioning. This impossibility relies on algebraic structures arising in incidence geometry over fields without proper subfields. Our argument connects recent advances in bounds on point-line incidences with previously developed techniques in algorithmic information theory and communication complexity. This suggests broader prospects for the algebraic and combinatorial methods developed in [20, 21, 22, 23, 19] for studying fundamental barriers in information theory and communication complexity. We emphasize that the natural Question 1 on page 8 (see also [13]) remains widely open.
Appendix A Halving the complexities of two strings
A.1 Halving complexities with a hashing argument
Proposition 3 (An. Muchnik,[31]).
Let $x$ and $y$ be arbitrary strings of length at most $n$. Then there exists a string $r$ of length $C(x|y)$ such that
  • $C(r|x)\stackrel{{\scriptstyle\log}}{{=}}0$ and
  • $C(x|r,y)\stackrel{{\scriptstyle\log}}{{=}}0$.
Corollary 2.
Let $x$ and $y$ be arbitrary strings of length at most $n$ and let $k$ be a number less than $C(x)$. Then there exists a string $r$ of length $k$ such that
  • $C(r)\stackrel{{\scriptstyle\log}}{{=}}k$,
  • $C(r|x)\stackrel{{\scriptstyle\log}}{{=}}0$, and
  • $I(r:y)\stackrel{{\scriptstyle\log}}{{=}}\max\{0,k-C(x|y)\}$.
Proof of the corollary.
First of all, we apply Proposition 3 and obtain a string $r^{\prime}$ of length $C(x|y)$ such that $C(r^{\prime}|x)\stackrel{{\scriptstyle\log}}{{=}}0$ and $C(x|r^{\prime},y)\stackrel{{\scriptstyle\log}}{{=}}0$. The latter condition means that the difference between $C(x|y)$ and $C(x|r^{\prime},y)$ is equal to $C(x|y)=C(r^{\prime})$. Therefore, $I(r^{\prime}:y)\stackrel{{\scriptstyle\log}}{{=}}0.$ If $k\leq C(x|y)$, we define $r$ as the prefix of $r^{\prime}$ of length $k$, and we are done. Otherwise, we apply again Proposition 3 and obtain $r^{\prime\prime}$ of length $C(x)-C(x|y)$ such that $C(r^{\prime\prime}|x)\stackrel{{\scriptstyle\log}}{{=}}0$ and $C(x|r^{\prime},r^{\prime\prime})\stackrel{{\scriptstyle\log}}{{=}}0$. Let $r^{\prime\prime\prime}$ be the prefix of $r^{\prime\prime}$ of length $k-C(x|y)$. We set $r=\langle r^{\prime},r^{\prime\prime\prime}\rangle$ (encoding of the pair of strings). It is easy to see that
$C(r)\stackrel{{\scriptstyle\log}}{{=}}C(r)+C(r^{\prime})\stackrel{{\scriptstyle\log}}{{=}}k$
and
$C(r|y)\stackrel{{\scriptstyle\log}}{{=}}C(r^{\prime}|y)\stackrel{{\scriptstyle\log}}{{=}}C(x|y),$
which implies $I(r:y)\stackrel{{\scriptstyle\log}}{{=}}k-C(x|y)$. ∎
Proof of Theorem 1.
Let us denote
$\alpha=C(a|b),\ \beta=C(b|a),\ \gamma=I(a:b),$
so that
$C(a)\stackrel{{\scriptstyle\log}}{{=}}\alpha+\gamma,\ C(b)\stackrel{{\scriptstyle\log}}{{=}}\beta+\gamma.$
We assume without loss of generality that $\alpha\leq\beta$.
We will construct $z$ as a combination of two components, $p$ and $q$, where $p$ and $q$ are suitable “hashes” of $a$ and $b$ respectively. To define the first of these components, we apply Corollary 2 and obtain $p$ such that
  • $C(p)=\frac{\alpha+\gamma}{2}$,
  • $C(p|a)\stackrel{{\scriptstyle\log}}{{=}}0$ (and, therefore, $I(p:a)=C(p)=\frac{\alpha+\gamma}{2}$),
  • $C(p|b)\stackrel{{\scriptstyle\log}}{{=}}C(p)$ (in other words, $I(p:b)\stackrel{{\scriptstyle\log}}{{=}}0$).
Case 1: $\alpha\geq\gamma$. Observe that in this case $\frac{\beta+\gamma}{2}<\beta$. We apply again Corollary 2 and take $q$ such that
  • $C(q)\stackrel{{\scriptstyle\log}}{{=}}\frac{\beta+\gamma}{2}$,
  • $C(q|b)\stackrel{{\scriptstyle\log}}{{=}}O(\log n)$ (and, therefore, $I(q:b)\stackrel{{\scriptstyle\log}}{{=}}C(q)\stackrel{{\scriptstyle\log}}{{=}}\frac{\beta+\gamma}{2}$),
  • $C(q|a)\stackrel{{\scriptstyle\log}}{{=}}C(q)$ (thus, $I(q:p)\stackrel{{\scriptstyle\log}}{{\leq}}I(q:a)\stackrel{{\scriptstyle\log}}{{=}}0$).
We let $z=\langle p,q\rangle$ (encoding of the pair ofstrings). A straightforward computation gives
$\begin{array}[]{rcl}C(a|z)&\stackrel{{\scriptstyle\log}}{{=}}&C(a|pq)\stackrel{{\scriptstyle\log}}{{=}}C(a,p,q)-C(p,q)\\ &\stackrel{{\scriptstyle\log}}{{=}}&C(a,q)-C(p,q)\stackrel{{\scriptstyle\log}}{{=}}C(a)+C(q)-C(p)-C(q)\\ &\stackrel{{\scriptstyle\log}}{{=}}&C(a)-C(p)\stackrel{{\scriptstyle\log}}{{=}}\alpha+\gamma-\frac{\alpha+\gamma}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{\alpha+\gamma}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{1}{2}C(a).\end{array}$
$\begin{array}[]{rcl}C(b|z)&\stackrel{{\scriptstyle\log}}{{=}}&C(b|pq)\stackrel{{\scriptstyle\log}}{{=}}C(b,p,q)-C(p,q)\\ &\stackrel{{\scriptstyle\log}}{{=}}&C(b,p)-C(p,q)\stackrel{{\scriptstyle\log}}{{=}}C(b)+C(p)-C(p)-C(q)\\ &\stackrel{{\scriptstyle\log}}{{=}}&C(b)-C(q)\stackrel{{\scriptstyle\log}}{{=}}\beta+\gamma-\frac{\beta+\gamma}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{\beta+\gamma}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{1}{2}C(b).\end{array}$
Case 2: $\alpha<\gamma$. In this case $p$ and $y$ are not independent, Corollary 2 gives
$I(p:y)\stackrel{{\scriptstyle\log}}{{=}}C(p)-C(a|b)\stackrel{{\scriptstyle\log}}{{=}}\frac{\alpha+\gamma}{2}-\alpha\stackrel{{\scriptstyle\log}}{{=}}\frac{\gamma-\alpha}{2}.$
It is easy to verify that $C(y|p)>\frac{\alpha+\beta}{2}.$ We apply one more time Corollary 2 and take $q$ such that
  • $C(q)\stackrel{{\scriptstyle\log}}{{=}}\frac{\alpha+\beta}{2}$,
  • $C(q|b)\stackrel{{\scriptstyle\log}}{{=}}0$ (and, therefore, $I(q:b)\stackrel{{\scriptstyle\log}}{{=}}C(q)\stackrel{{\scriptstyle\log}}{{=}}\frac{\alpha+\beta}{2}$),
  • $C(q|a)\stackrel{{\scriptstyle\log}}{{=}}C(q)$ (thus, $I(q:p)\stackrel{{\scriptstyle\log}}{{\leq}}I(q:a)\stackrel{{\scriptstyle\log}}{{=}}0$).
Similarly to Case 1 we let $z=\langle p,q\rangle$ and obtain the necessary conditional complexities: the computation
$\begin{array}[]{rcl}C(a|z)\stackrel{{\scriptstyle\log}}{{=}}\frac{\alpha+\gamma}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{1}{2}C(a).\end{array}$
is exactly the same as in Case 1, and
$\begin{array}[]{rcl}C(b|z)&\stackrel{{\scriptstyle\log}}{{=}}&C(b|pq)\stackrel{{\scriptstyle\log}}{{=}}C(b,p,q)-C(p,q)\\ &\stackrel{{\scriptstyle\log}}{{=}}&C(b,p)-C(p,q)\stackrel{{\scriptstyle\log}}{{=}}C(b|p)+C(p)-C(p)-C(q)\\ &\stackrel{{\scriptstyle\log}}{{=}}&C(b)-I(p:b)-C(q)\stackrel{{\scriptstyle\log}}{{=}}\beta+\gamma-\frac{\gamma-\alpha}{2}-\frac{\alpha+\beta}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{\beta+\gamma}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{1}{2}C(b).\end{array}$
A.2 Halving complexities with a topological argument
In this section, we presented an alternative (topological) proof for a special case of Theorem 1. This is the special case to which we repeatedly return as a running example throughout the paper: it involves a pair $(x,y)$ with the complexity profile
$C(x)\approx C(y)\approx 2n\quad\text{and}\quad I(x:y)\approx n.$
The argument outlined below can be combined with Muchnik’s hashing method from the previous section and extended to a significantly more general case; see [18].
Proposition 4.
Let $x,y$ be strings such that
$C(x)=2n+o(n),\ C(y)=2n+o(n),\ I(x:y)=n+o(n).$
Then there exists a string $z$ such that $C(x|z)\stackrel{{\scriptstyle\log}}{{=}}n\text{ and }C(y|z)\stackrel{{\scriptstyle\log}}{{=}}n.$
Proof.
(The argument below is a simplified version of the more general proof presented in [18, Theorem 4].) We will show that the required $z$ can be constructed as a pair $a=\langle x^{\prime},y^{\prime}\rangle$, where $x^{\prime}$ and $y^{\prime}$ are prefixes of $x$ and $y$ respectively. We are going to prove that such prefixes of $x$ and $y$ can be chosen properly, although we cannot specify their lengths explicitly
The construction of $z$ depends on two parameters: the length $\alpha$ of $x^{\prime}$ (which is an integer between 00 and $|x|$) and the length $\beta$ of $y^{\prime}$ (an integer between 00 and $|y|$). The possible values of these parameters can be thought of as a two-dimensional grid of points with integer coordinates, see the integer points inside the rectangle on the left in Fig. 3. For each integer pair $(\alpha,\beta)$ in this rectangle we get the corresponding prefixes $x^{\prime}=x_{[1:\alpha]}$ and $y^{\prime}=y_{[1:\beta]}$, and $z=\langle x^{\prime},y^{\prime}\rangle$. This $z$ gives us in turn a pair of values $(C(x|z),C(y|z)),$ which is a point in the picture on the right in Fig. 3. Observe that the mapping
$(\alpha,\beta)\mapsto(C(x\,|\,\langle x_{[1:\alpha]},y_{[1:\beta]}\rangle),C(y\,|\,\langle x_{[1:\alpha]},y_{[1:\beta]}\rangle)$ (13)
respects the Lipschitz condition: increasing $\alpha$ or $\beta$ by $1$, we add one bit to $x^{\prime}$ or $y^{\prime}$; this operation changes the values of $C(x|\langle x^{\prime},y\rangle)$ and $C(y|\langle x^{\prime},y\rangle)$ by only a bounded additive term $O(1)$. In the rest of the proof, our aim is to show that there exists a pair $(\alpha,\beta)$ on the left in Fig. 3 that is mapped to the $O(1)$-neighborhood of the point $(n,n)$ on the right.
$\alpha$$\beta$$(0,0)$$(|x|,0)$$(0,|y|)$$(|x|,|y|)$
$C(x|z)$$C(y|z)$$(0,0)$$(0,2n)$$(2n,0)$$(2n,2n)$$(n,n)$$(n,0)$$(0,n)$
Figure 3: Domain and codomain of the mapping (13)
First of, we observe that the points $(0,0)$, $(|x|,0)$, $(0,|y|)$, $(|x|,|y|)$ are mapped to the ($o(n)$-neighborhoods of) points
$(2n,2n),\ (0,n),\ (n,0),\ (0,0)$
on the right [^3] For example, $(\alpha,\beta)=(|x|,0)$ means that $z$ is a pair consisting of the entire $x$ and the empty prefix of $y$; conditional on this $z$, complexity of $x$ vanishes, and conditional complexity of $y$ shrinks to $C(y|x)=n\pm o(n)$. The other three cases can be settled in a similar way..
When we go along the vertical line $(|x|,0)$$(|x|,|y|)$, the string $z$ consists of the entire $x$ and a gradually growing prefix of $y$. Accordingly, the value of $C(x|z)$ remains equal to $O(1)$, and $C(y|z)$ decreases from $C(y|z)\pm O(\log n)$ to 00. That is, the image of $(\alpha,\beta)$ (taken from the left picture) runs through the segment $(0,n)$$(0,0)$ (on the right picture). Similarly, when $(\alpha,\beta)$ run the segment $(0,|y|)$$(|x|,|y|)$ on the left, the images run through the segment $(n,0)$$(0,0)$ on the right, at least within $o(n)$ precision.
Now let $(\alpha,\beta)$ runs through the segment $(0,0)$$(|x|,0)$. We know that the image of $(\alpha,\beta)$ (i.e., the corresponding values of $(C(x|z),C(y|z))$) gradually changes from $(2n,2n)$ to $(0,n)$. However, we cannot control exactly the behavior of $(C(x|z),C(y|z))$ on this interval: the image of $(0,0)$$(|x|,0)$ does not need to be a straight line; we only know that it is slowly changing line that descends from $(2n,2n)$ to $(0,n)$ and remains all the way above the horizontal line $C(y|z)=n-o(n)$ (indeed, the complexity value $C(y|x^{\prime})$ cannot be be much smaller than $C(y|x)=n-o(n)$).
Similarly, when $(\alpha,\beta)$ runs through the segment $(0,0)$$(0,|y|)$, the pair of values $(C(x|z),C(y|z))$ goes from $(2n,2n)$ to $(n,0)$. Thus, the border line of the rectangle on the left in Fig. 3 is mapped to the bold line on the right in the same figure. Observe that the point $(n,n)$ is inside (at least within $o(n)$ precision) the loop shown in bold on the right in Fig. 3. It remains to show that some point from the domain of the mapping (13) is mapped to the $O(1)$-neighborhood of $(n,n)$. In what follows we do it in three steps.
Step 1: From discrete mapping to a continuous function. The mapping in (13) is defined only on the points with integer coordinates. We can extend it to the points with real coordinates (still inside of the rectangle $0\leq\alpha\leq|x|$, $0\leq\beta\leq|y|$. To this end, we split the grid of integer points into triangles (as shown in Fig. 3); inside of each triangle, we define the function as the barycentric average of the values assigned to the vertices of this triangle. The extended function is continuous and even Lipschitz continuous (as the original discrete function respected the condition of Lipschitz).
Step 2: Exact real-valued preimage of the point $(n,n)$. In pour setting, we have a continuous mapping from a subset of $\mathbb{R}^{2}$ (the rectangle $0\leq\alpha\leq|x|$, $0\leq\beta\leq|y|$, which is homeomorphic to the closed disk) to $\mathbb{R}^{2}$. We know that borderline of the pre-image is mapped to a curve (the bold line on the right in Fig. 3) that goes around the point $(n,n)$. It follows immediately that some point $(\alpha_{0},\beta_{0})$ from the pre-image (possibly with non-integer coordinates) is mapped exactly to $(n,n)$. This fact is a basic result in algebraic topology, which is sometimes called the drum theorem [30] (informally, it claims that when we stretch an elastic film over a circle, we obtain a drum without holes in the middle; more formally, this fact is equivalent to the theorem stating that the circle is not a retract of the closed disk).
Step 3: Approximate integer pre-image of the point $(n,n)$. Since the mapping is Lipschitz continuous, we can replace $(\alpha_{0},\beta_{0})$ by the closest points $(\alpha_{0}^{\prime},\beta_{0}^{\prime})$ which is mapped to (13) to the $O(1)$-neighborhood of $(n,n)$. This observation concludes the proof of the theorem. ∎